### v2

# Motion-Aware Animatable Gaussian Avatars Deblurring

Abstract
The creation of 3D human avatars from multi-view videos is a significant yet challenging task in computer vision. However, existing techniques rely on high-quality, sharp images as input, which are often impractical to obtain in real-world scenarios due to variations in human motion speed and intensity. This paper introduces a novel method for directly reconstructing sharp 3D human Gaussian avatars from blurry videos. The proposed approach incorporates a 3D-aware, physics-based model of blur formation caused by human motion, together with a 3D human motion model designed to resolve ambiguities in motion-induced blur. This framework enables the joint optimization of the avatar representation and motion parameters from a coarse initialization. Comprehensive benchmarks are established using both a synthetic dataset and a real-world dataset captured with a 360-degree synchronous hybrid-exposure camera system. Extensive evaluations demonstrate the effectiveness and robustness of the model across diverse conditions.
11footnotetext: Corresponding author.
1 Introduction
Motion blur arises when scene changes occur during camera exposure, degrading image quality and perceptual clarity. Despite advances in camera technology, it remains prevalent due to the unpredictable nature of object motion. Consequently, many methods have been proposed to restore sharp details from blurry captures [63, 69, 79, 3, 87, 90].
In the field of 3D reconstruction, creating high-quality 3D human avatars holds immense potential for the industry [19, 68, 14, 35, 4, 15, 39, 84, 75, 5, 50, 81]. Cutting-edge methods employing 3D Gaussian Splatting [22, 9, 41, 10] and the Skinned Multi-Person Linear (SMPL) model [42] have achieved notable success. These techniques [14, 57, 13, 35] typically rely on video data with SMPL parameters from calibration techniques like EasyMocap [7, 56].
Figure 1: The ambiguity brought by motion blur. When reconstructing sharp 3DGS avatars from blurry frames, motion-induced blur introduces challenging ambiguities in motion interpretation.
While these models have achieved remarkable results with well-calibrated and high-quality video frames, their performance can be significantly degraded by blur effects caused by human motion. In practical scenarios, such motion-induced blur is often unavoidable due to the unpredictable speeds and movements of subjects. In particular, blur effects can adversely affect the performance of existing models in two key ways. First, blurry captures may cause the 3DGS model to learn a distorted 3D representation, arising from the inherent ambiguity of motion blur (Fig. 1). Such ambiguity hinders the accurate recovery of structural information and texture details. Second, even though static cameras can be calibrated before recording, blurry captures still lead to erroneous estimation of SMPL parameters. [^1] Some studies, such as [14], incorporate a correction module to refine pose parameters during training. However, experimental results indicate that these modifications often lead to suboptimal performance due to the absence of a precise blur-aware motion modeling process.
A practical strategy to mitigate this issue is to implement a two-stage baseline approach. First, 2D deblurring techniques [80, 86, 53, 36] are applied to restore sharp video sequences. Then, the deblurred frames are used to train the 3DGS avatar model. While this baseline method improves visual quality by partially resolving motion blur ambiguity through 2D deblurring, it overlooks intrinsic 3D scene information. This omission often leads to inconsistencies across multiple views during the deblurring process, compromising the performance of the 3DGS avatar model.
Based on the settings of existing gaussian avatar models [14, 57, 13], this paper introduces the first model for reconstructing sharp, animatable 3D human avatars directly from blurry video frames. By exploiting the distinctive characteristics of human avatar models, we expand the conventional physics-based 2D image blur process to a 3D-aware blur formation model. This formulation decomposes the inherently ill-posed deblurring problem into two primary tasks: optimizing sub-frame motion representations and constructing the canonical sharp 3DGS avatar model. To mitigate the sub-frame motion ambiguity introduced by motion-induced blur, a 3D-aware human motion model based on the SMPL framework is incorporated. This model enables the joint optimization of human motion and the sharp 3DGS avatar. Using the optimized motion and canonical 3DGS, motion-blurred frames are synthesized by averaging a sequence of “virtual” sharp images, and the loss is computed against the observed blurry frames.
Extensive evaluations are conducted on various datasets, baselines, and settings to confirm the superiority and robustness of the proposed method. As there are no blur-aware avatar benchmarks for qualitative and quantitative evaluation, a synthetic dataset is developed based on the widely used ZJU-MoCap dataset [56], complemented by a high-quality real-world dataset captured using a 360-degree hybrid-exposure system. A do-it-yourself style demo with iPhone 16 Pro is also provided. Codes and datasets will be publicly released to facilitate relevant research in the future.
2 Related Work
Video Deblurring. In recent years, deblurring has shifted from traditional methods [18, 31, 59, 73] to learning-based models that directly predict sharp videos [48, 65, 89, 17, 49, 71, 88, 91, 6, 92, 52, 60, 27, 28, 37, 2]. Recently, the advent of NeRF [46] and 3DGS [22] has promoted multi-view deblurring. Specifically, methods like [44, 67, 29, 85] focus on deblurring static scenes affected by defocus or camera movement, while some approaches [43, 64, 1, 61] deblurs dynamic scenes. However, none of these methods could reconstruct sharp and animatable avatars from blurry videos.
3D Human Avatars. Since the advent of Neural Radiance Fields [46], research on neural human rendering has flourished [51, 62, 74, 12, 20, 21, 34, 54, 55, 68, 72, 78]. Recently, the superior performance of 3DGS [22] has spurred extensive research into using 3D Gaussian representations for dynamic human reconstruction [93, 35, 19, 47, 76, 26, 30, 40, 13]. However, these methods rely on high-quality input images and fail when applied to blurry captures.
3 Method
3.1 3D Blur Formation Model
The physical mechanism of image formation involves capturing photons during the camera’s exposure period. This process can be mathematically represented in the 2D camera coordinate space as the integration of a sequence of conceptual sharp images over the exposure duration:
$\mathbf{I}^{B}(\mathbf{u})=\int^{\tau}_{0}\mathbf{I}^{S}_{t}(\mathbf{u})\,\mathrm{d}t,$ (1)
where $\mathbf{I}^{B}(\mathbf{u})\in\mathbb{R}^{3\times H\times W}$ denotes the captured blurry image, with $H$ and $W$ representing the image height and width, respectively. The variable $\mathbf{u}\in\mathbb{R}^{2}$ indicates the pixel location, $\tau$ is the exposure time, and $\mathbf{I}^{S}_{t}(\mathbf{u})\in\mathbb{R}^{3\times H\times W}$ represents the virtual sharp image at time $t$ within the exposure period. This continuous integration process can be discretely approximated as:
$\mathbf{I}^{B}(\mathbf{u})\approx\frac{1}{n}\sum^{n-1}_{i=0}\mathbf{I}^{S}_{i}(\mathbf{u}),$ (2)
where the blurred image $\mathbf{I}^{B}(\mathbf{u})$ is estimated by averaging $n$ virtual intermediate sharp images $\mathbf{I}^{S}_{i}(\mathbf{u})$. The blur formation process is redefined from the perspective of 3D human avatar modeling, extending beyond the constraints of the 2D camera coordinate space. Specifically, the motion of a 3DGS avatar during the exposure period is modeled using a set of $K$ 3D Gaussians ${G_{k}(\mathbf{x})}_{k=0}^{K-1}$ defined in the canonical space, together with a sequence of $T$ SMPL [42] parameters $\{\mathcal{S}_{t}\}^{T-1}_{t=0}=\{\bm{\Theta}_{t},\bm{\beta}_{t},\mathcal{B}_{t}\}^{T-1}_{t=0}$. These parameters dynamically deform the 3D Gaussians into the observation space at each discrete time step $t$. Consequently, the resulting blurry image $\mathbf{I}^{B}$ can be expressed as:
$\mathbf{I}^{B}=\frac{1}{T}\sum^{T-1}_{t=0}\mathcal{R}(\mathcal{W}(\{G_{k}(\mathbf{x})\}^{K-1}_{k=0},\mathcal{S}_{t}),\bm{\mathrm{R}},\bm{\mathrm{K}}),$ (3)
where $\mathbf{x}$ denotes the coordinates of the 3D Gaussians, and $\bm{\mathrm{R}}$ and $\bm{\mathrm{K}}$ represent the camera’s extrinsic and intrinsic parameters, respectively. The operator $\mathcal{R}$ denotes the rasterization process of the 3DGS model, while $\mathcal{W}$ represents the warping of 3D Gaussians from the canonical space to the observation space, governed by the SMPL parameters $\mathcal{S}_{t}$.
3.2 3D Human Motion Model
As illustrated in Fig. 1, motion-induced blur introduces significant ambiguities in motion that are readily observable in the pixel space. To tackle this challenge, a 3D-aware human motion model is proposed to effectively estimate both sub-frame motion within each exposure period and inter-frame global motion across consecutive frames.
Sub-frame rigid sequential pose model. The pose parameters $\bm{\Theta}_{t}$, defined in the SMPL model as $\bm{\Theta}_{t}\in\mathbb{R}^{24\times 3}$, represent the rotations of $24$ joints at each time step $t$, expressed in $\textbf{SO}(3)$. Considering the inherent continuity of joint motion, De Boor–Cox formulation for B-splines [58, 8, 66, 33] is employed to interpolate intermediate poses.
For each joint $j$ in the SMPL model, $P$ control parameters are defined as $\tilde{\bm{\Theta}}^{j}=\{\tilde{\bm{\Theta}}^{j}_{p}\}^{P-1}_{p=0}\in\mathbb{R}^{P\times 3}$, where $P$ denotes the predefined number of control knots. These learnable parameters are initialized from the coarse estimation, and optimized during the training process. At each time step $t$, the interpolation process begins with the computation of the timestep basis $\mathbf{B}(t)\in\mathbb{R}^{1\times P}$:
$\mathbf{B}(t)=[1,\frac{t}{T},(\frac{t}{T})^{2},\ldots,(\frac{t}{T})^{P-1}],$ (4)
where $T$ denotes the total exposure time. Based on this basis, the interpolated pose parameters for the $j$-th joint at time step $t$ are computed as:
$\hat{\bm{\Theta}}^{j}_{t}=\mathbf{B}(t)\cdot\mathcal{M}^{P}\cdot\tilde{\bm{\Theta}}^{j},$ (5)
where $\mathcal{M}^{P}\in\mathbb{R}^{P\times P}$ is the interpolation matrix [58]:
$\mathcal{M}^{P}_{i,j}=C^{P-1}_{P-1-i}\sum_{s=j}^{P-1}(-1)^{s-j}C^{P}_{s-j}(P-s-1)^{P-1-i},$ (6)
where $C^{n}_{k}=\frac{n!}{k!(n-k)!}$ denotes the binomial coefficient.
Figure 2: Brief illustration of the pipeline. The sub-frame motion for each blurry frame is modeled using the SMPL representation, followed by warping the canonical 3DGS according to the estimated motion parameters. The final blurry image is synthesized by averaging the sequence of rendered virtual sharp images.
Figure 3: 360-degree hybrid-exposure camera system. Left: the inner side and outer side of the capture cage. Right: illustration of the system and samples of captures.
Capture type Blur Sharp
View number 44 88
Exposure $50\text{\,}\mathrm{m}\mathrm{s}$ $3.125\text{\,}\mathrm{m}\mathrm{s}$
Resolution $2448\times 2048$
Camera model BFS-U3-51S5C
Scene number 88
Picture format PNG
Table 1: Configuration of the captured real data.
Figure 4: Qualitative comparison results on synthetic dataset. Zoom in for the best view.
Pose deformation model. While B-spline interpolation captures basic pose trajectories, it remains limited in representing non-rigid, high-frequency pose variations. To overcome this, a pose displacement $\Delta^{j}_{t}$ is introduced for each joint $j$ at every time step $t$ within the exposure period:
$\Delta^{j}_{t}$ $=G_{disp}(\hat{\bm{\Theta}}^{j}_{t};\theta_{disp}),$ (7)
$\bm{\Theta}^{j}_{t}$ $=\hat{\bm{\Theta}}^{j}_{t}+\Delta^{j}_{t},$ (8)
where $G_{disp}$ denotes the CNN designed to estimate fine-grained pose variations, and $\theta_{disp}$ represents its set of learnable parameters. This formulation allows the model to more accurately capture complex pose dynamics, restoring more realistic motion within each blurred frame.
Inter-frame motion regularization. Although the L1 loss can yield satisfactory photometric results at the midpoint of the exposure period $(t=0.5)$, it may suffer from directional ambiguity, as illustrated in Fig. 1(c). This ambiguity occurs because motion in either direction can produce a similarly plausible blurry image, potentially leading to inaccurate motion direction estimation and misaligned renderings at non-midpoint timesteps. To mitigate this issue, a regularization term is introduced based on the inter-frame continuity commonly observed in video sequences. This term measures the Geodesic distance between the pose parameters at the final timestep of the current exposure period and those at the initial timestep of the subsequent exposure period:
$\mathcal{L}_{reg}=\frac{1}{24\cdot(N_{e}-1)}\sum^{N_{e}-2}_{n=0}\sum^{23}_{j=0}\left|\hat{\mathbf{\Theta}}^{j}_{n,T-1}-\hat{\mathbf{\Theta}}^{j}_{n+1,0}\right|_{\text{G}},$ (9)
where $\hat{\mathbf{\Theta}}^{j}_{n,t}$ denotes the estimated pose parameters of the $j$-th joint at timestep $t$ for the $n$-th exposure period, and $N_{e}$ represents the total number of exposure periods. $\left|\cdot\right|_{\text{G}}$ denotes the Geodesic distance. This regularization term is designed to enhance the inter-frame temporal coherence of joint movements, ensuring that the motion remains consistent and progresses naturally from one frame to the next.
Shape estimation. The shape parameters $\hat{\bm{\beta}}\in\mathbb{R}^{10}$ of the SMPL model are initialized from the coarse estimation, then optimized as parameters during the training process.
[…]
where $\delta$ is predicted by a simple CNN that takes the coordinates of all 3D Gaussians as input, enabling dynamic refinement of the LBS weights.
3.3 Optimization
Pipeline. A brief illustration for the optimization pipeline of the proposed model is provided in Fig. 2. The process begins with the estimation of sub-frame motion using the dedicated motion model described in Sec. 3.2. The SMPL parameters are initialized using the coarse estimation from blurry frames, and gradually optimized within the 3D-aware framework. The canonical 3D Gaussian Splatting (3DGS) representation is then transformed according to the estimated motion parameters. Subsequently, sharp virtual images are rendered for each intermediate timestep. These images are averaged to produce the final blurry image used for loss computation.
Figure 5: Time synchronization of the camera system. ‘TD’ and ‘EX’ stand for “Trigger Delay” and “Exposure”.
Loss function. The loss function includes the L1 loss between the synthesized blurry frame $\hat{\mathbf{I}}^{B}$ and the observed blurry frame $\mathbf{I}^{B}$, and the inter-frame motion regularization loss as decribed in Sec. 3.2:
[…]
The total loss enforces both accurate avatar reconstruction and the estimation of realistic, temporally coherent motion.
4 Experiments
Figure 6: Qualitative comparison results on real data. Zoom in for the best view.
Figure 7: Qualitative results for ablation studies. Zoom in for the best view.
Table 2: Quantitative comparison results on two datasets. We colorize results as best, second best, and third best.
Methods Synthetic Dataset Real Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
GauHuman 23.080 0.7660 0.2277 25.602 0.8044 0.2380
ShiftNet + GauHuman 23.089 0.7695 0.2219 25.549 0.8043 0.2347
RVRT + GauHuman 23.078 0.7697 0.2218 25.547 0.8065 0.2343
VRT + GauHuman 23.074 0.7696 0.2205 25.553 0.8067 0.2345
BSST + GauHuman 23.081 0.7698 0.2212 25.568 0.8068 0.2342
Ours 25.546 0.8290 0.1476 27.010 0.8271 0.1668
4.1 Benchmarks
Synthetic dataset. The synthetic data is constructed based on the ZJU-MoCap [56] dataset since it is widely adopted by many 3D human avatar research work [14, 11, 72, 57, 23, 77]. Following their practice, six human subjects (IDs: $377$, $386$, $387$, $392$, $393$, $394$) are selected for experiments. Each blurry frame was synthesized by averaging $K_{blur}$ consecutive sharp frames To mitigate discontinuities caused by direct averaging, the RIFE model [16] is employed to interpolate $16$ intermediate frames between adjacent sharp frames. Blurry frames from Cameras $04$, $10$, $16$, and $22$ are used for training, while the sharp frames from the remaining $18$ cameras are reserved for novel-view test. SMPL parameters are calibrated using the blurry video with EasyMoCap [7, 56], and human masks are obtained using SAM [25].
Table 3: Quantitative ablation results on two datasets. We colorize results as best, second best, and third best.
Models Synthetic Dataset Real Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o interp. 24.009 0.8053 0.1620 25.825 0.8140 0.1729
w/o pose deform 25.301 0.8229 0.1545 26.426 0.8184 0.1743
w/o LBS opt. 25.394 0.8261 0.1486 26.821 0.8233 0.1697
w/o shape opt. 25.529 0.8284 0.1481 26.964 0.8261 0.1669
Ours (full model) 25.546 0.8290 0.1476 27.010 0.8271 0.1668
Figure 8: Qualitative ablation results for $\mathcal{L}_{reg}$. Model w/o $\mathcal{L}_{reg}$ gives worse result for non-middle timesteps due to incorrect motion direction estimates. Error maps are visualized for each output.
Real dataset. The dataset consists of real-captured blurry videos and synchronously captured sharp videos from novel viewpoints. The capture system employs $12$ time-synchronized Blackfly BFS-U3-51S5C cameras arranged in a 360-degree spatial configuration, as illustrated in Fig. 3. $4$ cameras record blurry videos, while the remaining $8$ capture sharp videos for novel-view evaluation. Fig. 5 illustrates the time synchronization of two types of cameras. To ensure consistent pixel intensity, the blurry cameras are equipped with a 1/16{1}/{16} ND filter. Detailed configuration are reported in Tab. 1. SMPL parameters are calibrated using EasyMoCap [7], and human masks are generated using SAM [25].
Table 4: Quantitative ablation results for the inter-frame motion regularization loss $\mathcal{L}_{reg}$. We colorize better result of each sector.
$K_{blur}$ Models Middle Timestep Non-middle Timesteps
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
55 w/o $\mathcal{L}_{reg}$ 25.567 0.8296 0.1478 24.421 0.8111 0.1601
w/ $\mathcal{L}_{reg}$ 25.546 0.8290 0.1476 25.417 0.8269 0.1485
77 w/o $\mathcal{L}_{reg}$ 25.113 0.8197 0.1585 23.737 0.7950 0.1758
w/ $\mathcal{L}_{reg}$ 25.155 0.8200 0.1557 25.036 0.8179 0.1567
99 w/o $\mathcal{L}_{reg}$ 24.628 0.8114 0.1680 23.198 0.7825 0.1888
w/ $\mathcal{L}_{reg}$ 24.680 0.8126 0.1636 24.605 0.8111 0.1645
$11$ w/o $\mathcal{L}_{reg}$ 24.241 0.8009 0.1786 22.926 0.7717 0.1998
w/ $\mathcal{L}_{reg}$ 24.353 0.8039 0.1725 24.317 0.8031 0.1726
Implementation details. Adam optimizer [24] is used with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rates and decay schedules follow the original 3DGS [22]. Following the protocol in [14], the input resolution is set to $512\times 512$ for the synthetic dataset and $612\times 512$ for the real dataset. All models are trained on a single NVIDIA GeForce RTX 4090 GPU. Since the images contain large black regions, each image is cropped to the human bounding box for quantitative evaluation.
4.2 Comparisons
The proposed model is compared against several baseline approaches to assess its effectiveness. The simplest baseline involves training a SOTA avatar model (e.g., GauHuman [14]) using blurry video frames. A two-stage baseline is then considered, where 2D deblurring methods (RVRT [38], ShiftNet [32], VRT [36], and BSST [83]) are first applied to restore sharp frames, then used to train the avatar model.
Quantitative results. Quantitative comparisons on two datasets are presented in Tab. 2. The two-stage baselines exhibit suboptimal performance, with noticeably lower quantitative metrics. This limitation arises because 2D deblurring models fail to maintain multi-view consistency across different viewpoints and lack explicit modeling of the scene’s intrinsic 3D structure. They only slightly outperform the direct 3DGS avatar model, highlighting that inconsist 2D deblurring constrains the 3DGS model’s ability to produce accurate reconstructions. In contrast, the proposed method outperforms baselines by explicitly modeling the 3D motion-induced blur formation process.
Figure 9: Visualization of the initial estimated SMPL by EasyMocap and optimized SMPL sequence by the proposed model.
Table 5: Qualitative ablation results on trajectories representations. We colorize result as best, second best, and third best.
Models Synthetic Dataset Real Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Linear 25.516 0.8289 0.1483 26.899 0.8257 0.1680
Slerp 25.539 0.8288 0.1481 27.013 0.8264 0.1674
B-Spline 25.546 0.8290 0.1476 27.010 0.8271 0.1668
Table 6: Ablation results for control knot number $P$.
$P$ Synthetic Dataset Real Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
22 25.516 0.8289 0.1480 26.899 0.8257 0.1680
33 25.482 0.8284 0.1485 27.003 0.8266 0.1671
44 25.546 0.8290 0.1476 27.010 0.8271 0.1668
55 25.540 0.8288 0.1490 26.977 0.8260 0.1673
Qualitative results. Qualitative comparison results are shown in Fig. 4 and Fig. 6. While integrating 2D deblurring techniques with the 3DGS avatar model provides minor improvements in reconstruction quality (e.g., the arm region in Fig. 4), prominent visual artifacts remain, including residual blur and loss of detail around body contours. These artifacts arise from inconsistencies in deblurring across multiple viewpoints, as 2D deblurring methods do not enforce multi-view consistency. The proposed approach incorporates a physics-based blur formation model, allowing accurate learning of the 3D representation and enabling simultaneous deblurring and improved avatar reconstruction.
4.3 Ablation Study
Model components. Ablation studies are conducted to evaluate the contributions of different model components.
  • w/o interp.: Independently optimizing the pose at each timestep instead of using pose interpolation.
  • w/o pose deform: Using only the rigid sequential pose model, excluding pose deformation model.
  • w/o LBS opt.: Employing pre-trained LBS weights without additional optimization.
  • w/o shape opt.: Directly using calibrated SMPL shape parameters without further refinement.
The quantitative and qualitative results are presented in Tab. 3 and Fig. 7. Independently optimizing poses for each intermediate timestep leads to unordered motion estimations (the case illustrated in Fig. 1 (b)), leading to misaligned deblurring results. Excluding the pose deformation model introduces additional artifacts, as B-spline interpolation alone cannot adequately capture complex human motion. The performance also degrades when LBS weights and SMPL shape parameters are not refined, emphasizing the necessity of these components.
Figure 10: Visualizing SMPL under different perturbations.
Table 7: Ablation results for virtual sharp image number $T$.
$T$ Synthetic Dataset Real Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
33 25.528 0.8276 0.1503 26.837 0.8253 0.1695
55 25.546 0.8290 0.1476 27.010 0.8271 0.1668
77 25.532 0.8281 0.1499 26.993 0.8269 0.1670
99 25.522 0.8279 0.1501 26.965 0.8261 0.1676
Table 8: Quantitative results with different perturbations.
$\xi$ 0 $0.1$ $0.2$ $0.3$ $0.4$
PSNR 25.55 25.51 25.44 25.39 25.14
SSIM .8290 .8284 .8271 .8258 .8225
LPIPS .1476 .1476 .1493 .1504 .1524
Inter-frame regularization loss. Evaluations on $\mathcal{L}_{reg}$ are summarized in Tab. 4 and Fig. 8. It improves performance especially for non-middle timesteps by enabling more accurate estimations for inter-frame motion directions. Quantitative results on real data are excluded due to the lack of GT frames for non-middle timesteps.
Trajectory representations. Different trajectory representations are evaluated in the proposed model, including B-spline interpolation, linear interpolation, and spherical linear interpolation (Slerp). The corresponding results are presented in Tab. 5. It can be observed that B-spline interpolation generally achieves the best performance.
Hyperparameters. The proposed model is evaluated with varying numbers of control knots, $P=2,3,4,5$, for B-spline sequential pose interpolation. As reported in Tab. 6, the best performance is obtained with $P=4$. However, the performance differences remain relatively small, as the pose deformation model enhances robustness for smaller $P$ values by effectively modeling complex motions, thereby reducing performance gaps across different settings of $P$. The influence of virtual sharp image number $T$ is examined in Tab. 7. Optimal performance is achieved when $T=5$.
4.4 Robustness and Generality Evaluation
Figure 11: Visualization for different blur magnitudes $K_{blur}$.
Table 9: Quantitative result w/o optimizing SMPL parameters.
Methods Synthetic Dataset Real Dataset
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o SMPL opt. 21.63 .7269 .2421 25.14 .7855 .2423
Ours 25.55 .8290 .1476 27.01 .8271 .1668
Table 10: Quantitative comparisons with different $K_{blur}$.
$K_{blur}$ Metrics GauH ShiftNet RVRT VRT BSST Ours
+ GauH + GauH + GauH + GauH
77 PSNR $\uparrow$ 22.983 23.003 22.992 22.972 23.009 25.155
SSIM $\uparrow$ 0.7599 0.7648 0.7650 0.7654 0.7652 0.8200
LPIPS $\downarrow$ 0.2378 0.2314 0.2301 0.2299 0.2294 0.1557
99 PSNR $\uparrow$ 22.693 22.699 22.692 22.693 22.732 24.680
SSIM $\uparrow$ 0.7532 0.7585 0.7587 0.7588 0.7590 0.8126
LPIPS $\downarrow$ 0.2497 0.2426 0.2411 0.2412 0.2409 0.1636
$11$ PSNR $\uparrow$ 22.635 22.632 22.644 22.647 22.652 24.353
SSIM $\uparrow$ 0.7549 0.7554 0.7555 0.7564 0.7557 0.8039
LPIPS $\downarrow$ 0.2552 0.2502 0.2477 0.2480 0.2482 0.1725
Influence of the accuracy of SMPL estimation. SMPL estimations from blurry data could be inaccurate. However, the proposed approach does not depend on precise SMPL estimates but only coarse initializations. Fig. 9 visualizes the initial and optimized SMPL parameters, showing that the method refines inaccurate initializations and recover accurate sub-frame SMPL poses. To further assess robustness against inaccurate initialization, random perturbations $\epsilon\sim U[-\xi,\xi]$ are added to the initializations. Fig. 10 visualizes the effects of varying $\xi$ values. The quantitative results in Tab. 8 confirm the method’s resilience to even large perturbations. Finally, experiments are conducted without optimizing SMPL (pose, shape, LBS), using only the initial estimates. The results in Tab. 9 proves the necessity of SMPL optimization.
Different blur magnitudes $K_{blur}$. To evaluate the robustness of the proposed method under varying blur intensities, additional datasets were synthesized with blur magnitudes $K_{blur}=7,9,11$. Visualization of different blur magnitudes $K_{blur}$ is provided in Fig. 11. As reported in Tab. 10, the proposed method consistently outperforms all baseline approaches across varying blur level. Furthermore, corresponding ablation results in Tab. 11 confirm the contribution of each model component under different blur intensities.
Figure 12: Visualization of different mask parsing results.
Table 11: Quantitative ablation results for different $K_{blur}$.
$K_{blur}$ Metrics w/o w/o pose w/o w/o Ours
interp. deform LBS opt. shape opt. (full model)
77 PSNR $\uparrow$ 23.611 24.813 25.050 25.153 25.155
SSIM $\uparrow$ 0.7962 0.8127 0.8180 0.8199 0.8200
LPIPS $\downarrow$ 0.1710 0.1640 0.1563 0.1561 0.1557
99 PSNR $\uparrow$ 22.800 24.239 24.596 24.683 24.680
SSIM $\uparrow$ 0.7811 0.8035 0.8109 0.8127 0.8126
LPIPS $\downarrow$ 0.1850 0.1734 0.1638 0.1639 0.1636
$11$ PSNR $\uparrow$ 22.350 23.754 24.282 24.345 24.353
SSIM $\uparrow$ 0.7695 0.7924 0.8022 0.8037 0.8039
LPIPS $\downarrow$ 0.1945 0.1852 0.1725 0.1726 0.1725
Table 12: Quantitative results with different mask estimations.
Mask PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
sharp 25.61 .8309 .1465
SAM 25.55 .8290 .1476
rembg 25.49 .8281 .1488
Different view numbers $N_{view}$. To assess robustness with respect to the view numbers, the proposed model is evaluated using varying numbers of training views $N_{view}=2,3,4$. As shown in Tab. 13, the proposed method consistently outperforms baseline approaches with different $N_{view}$, demonstrating its strong generality. Furthermore, the ablation results under different $N_{view}$ in Tab. 14 further confirm the effectiveness of each component.
Mask estimation. Additional experiments are conducted to evaluate the model’s robustness to inaccurate masks due to blurriness. Since the synthetic dataset is generated from sharp frames, an “ideal” case is defined in which the masks are obtained by overlapping estimated masks from the corresponding sharp frames. Though such masks cannot be accessed in applications, by comparing to this idealized setting, the robustness of the model to mask estimation can be evaluated. Furthermore, alternative mask tools like rembg are adapted. As summarized in Tab. 12, the proposed method performs comparably to the ideal case, and the variation across different mask models remains minimal. Fig. 12 visualizes different mask estimations.
Figure 13: Novel pose animation results using out-of-domain motion from the AMASS [45] dataset.
Table 13: Quantitative comparisons with different $N_{view}$.
$N_{view}$ Metrics GauH ShiftNet RVRT VRT BSST Ours
+ GauH + GauH + GauH + GauH
2 PSNR $\uparrow$ 21.844 21.805 21.805 21.799 21.801 23.331
SSIM $\uparrow$ 0.7501 0.7533 0.7533 0.7534 0.7532 0.7712
LPIPS $\downarrow$ 0.2506 0.2436 0.2431 0.2435 0.2431 0.1891
3 PSNR $\uparrow$ 22.514 22.580 22.594 22.574 22.590 24.979
SSIM $\uparrow$ 0.7545 0.7584 0.7580 0.7580 0.7581 0.8134
LPIPS $\downarrow$ 0.2317 0.2225 0.2219 0.2225 0.2229 0.1587
4 PSNR $\uparrow$ 23.080 23.089 23.078 23.074 23.081 25.546
SSIM $\uparrow$ 0.7660 0.7695 0.7697 0.7696 0.7698 0.8290
LPIPS $\downarrow$ 0.2277 0.2219 0.2218 0.2205 0.2212 0.1476
Novel-pose evaluation. Novel poses from the AMASS dataset [45] are used to animate the reconstructed avatars for qualitative comparison. Fig. 13 shows that the proposed model achieves the best result.
4.5 Further Discussions
4.5.1 Demo captures from smartphones
To further demonstrate the generality of the proposed method in real-world scenarios, monocular video is captured using an iPhone 16 Pro, and SMPL parameters are estimated with TRAM [70], an open-sourced human motion estimation model. Fig. 14 confirm that the method generalizes well to real-world data and exhibits strong robustness to monocular SMPL estimations.
4.5.2 Limitations
While the proposed method successfully estimates motion and generates sharp 3DGS avatars, it has certain limitations. For example, since the model is built upon 3DGS representation, it struggles to accurately recover the geometry of the person, e.g., the surface normal or the BRDF. In addition, real-world blur occurs as light integrates in the linear radiance space before the ISP process. Averaging non-linear sRGB values directly can lead to inaccuracies, particularly in high-contrast regions where linear summation is essential for physical fidelity. Future work aims to explore these aspects. Also, the proposed model relies on SMPL to represent sub-frame motion for 3D-aware deblurring. However, when the subject has handheld objects or wears loose garments, it fails to recover the motion of these accessories because SMPL does not have corresponding “joints”. Future work could explore adaptive joints [82] with non-rigid modeling under blurriness to support more versatile avatar deblurring.
Figure 14: Qualitative results on captures from an iPhone 16 Pro.
Table 14: Quantitative ablation results with different $N_{view}$.
$N_{view}$ Metrics w/o w/o pose w/o w/o Ours
interp. deform LBS opt. shape opt.
2 PSNR $\uparrow$ 22.661 23.310 23.321 23.411 23.331
SSIM $\uparrow$ 0.7584 0.7704 0.7718 0.7737 0.7712
LPIPS $\downarrow$ 0.1978 0.1919 0.1896 0.1893 0.1891
3 PSNR $\uparrow$ 23.927 24.808 24.899 24.921 24.979
SSIM $\uparrow$ 0.7973 0.8081 0.8115 0.8129 0.8134
LPIPS $\downarrow$ 0.1683 0.1656 0.1594 0.1589 0.1587
4 PSNR $\uparrow$ 24.009 25.301 25.394 25.529 25.546
SSIM $\uparrow$ 0.8053 0.8229 0.8261 0.8284 0.8290
LPIPS $\downarrow$ 0.1620 0.1545 0.1486 0.1481 0.1476
5 Conclusion
In this paper, a novel approach is presented for reconstructing sharp 3D human avatars from blurry video frames. By extending the traditional 2D motion blur formulation into a 3D-aware blur formation model, the proposed method jointly optimizes sub-frame motion representations while learning a canonical 3DGS human avatar model. To establish a benchmark for this task, both a synthetic dataset derived from the ZJU-MoCap and a real-captured dataset are provided. Extensive evaluations demonstrate that the proposed model consistently outperforms existing baselines.

### v3

@@
 Comprehensive benchmarks are established using both a synthetic dataset and a real-world dataset captured with a 360-degree synchronous hybrid-exposure camera system.
-Extensive evaluations demonstrate the effectiveness and robustness of the model across diverse conditions.
+Extensive evaluations demonstrate the effectiveness of the model across diverse conditions.
+Codes Available: https://github.com/MyNiuuu/MAD-Avatar
 11footnotetext: Corresponding author.
@@
 Cutting-edge methods employing 3D Gaussian Splatting [21, 8, 40, 9] and the Skinned Multi-Person Linear (SMPL) model [41] have achieved notable success.
-These techniques [14, 57, 13, 35] typically rely on video data with SMPL parameters from calibration techniques like EasyMocap [7, 56].
+These techniques [13, 57, 12, 34] typically rely on video data with SMPL parameters from calibration techniques like EasyMocap [56].
 Figure 1: The ambiguity brought by motion blur. When reconstructing sharp 3DGS avatars from blurry frames, motion-induced blur introduces challenging ambiguities in motion interpretation.
@@
 Second, even though static cameras can be calibrated before recording, blurry captures still lead to erroneous estimation of SMPL parameters.
-[^1] Some studies, such as [14], incorporate a correction module to refine pose parameters during training.
-However, experimental results indicate that these modifications often lead to suboptimal performance due to the absence of a precise blur-aware motion modeling process.
+[^1] While certain approaches, such as GauHuman [13], integrate pose refinement modules, they often yield suboptimal results because they lack a rigorous mechanism for modeling blur-aware motion.
 A practical strategy to mitigate this issue is to implement a two-stage baseline approach.
@@
 The pose parameters $\bm{\Theta}_{t}$, defined in the SMPL model as $\bm{\Theta}_{t}\in\mathbb{R}^{24\times 3}$, represent the rotations of $24$ joints at each time step $t$, expressed in $\textbf{SO}(3)$.
-Considering the inherent continuity of joint motion, De Boor–Cox formulation for B-splines [58, 8, 66, 33] is employed to interpolate intermediate poses.
+Considering the inherent continuity of joint motion, De Boor–Cox formulation for B-splines [58, 7, 66, 32, 50] is employed to interpolate intermediate poses.
 For each joint $j$ in the SMPL model, $P$ control parameters are defined as $\tilde{\bm{\Theta}}^{j}=\{\tilde{\bm{\Theta}}^{j}_{p}\}^{P-1}_{p=0}\in\mathbb{R}^{P\times 3}$, where $P$ denotes the predefined number of control knots.
@@
 Figure 7: Qualitative results for ablation studies. Zoom in for the best view.
+4.1 Benchmarks
+Synthetic dataset.
+The synthetic data is constructed based on the ZJU-MoCap [56] dataset since it is widely adopted by many 3D human avatar research work [13, 10, 72, 57, 22, 77].
+Following their practice, six human subjects (IDs: $377$, $386$, $387$, $392$, $393$, $394$) are selected for experiments.
+Each blurry frame was synthesized by averaging $K_{blur}$ consecutive sharp frames To mitigate discontinuities caused by direct averaging, the RIFE model [15] is employed to interpolate $16$ intermediate frames between adjacent sharp frames.
+Blurry frames from Cameras $04$, $10$, $16$, and $22$ are used for training, while the sharp frames from the remaining $18$ cameras are reserved for novel-view test.
+SMPL parameters are calibrated using the blurry video with EasyMoCap [56], and human masks are obtained using SAM [24].
 Table 2: Quantitative comparison results on two datasets. We colorize results as best, second best, and third best.
@@
 Ours 25.546 0.8290 0.1476 27.010 0.8271 0.1668
-4.1 Benchmarks
-Synthetic dataset.
-The synthetic data is constructed based on the ZJU-MoCap [56] dataset since it is widely adopted by many 3D human avatar research work [14, 11, 72, 57, 23, 77].
-Following their practice, six human subjects (IDs: $377$, $386$, $387$, $392$, $393$, $394$) are selected for experiments.
-Each blurry frame was synthesized by averaging $K_{blur}$ consecutive sharp frames To mitigate discontinuities caused by direct averaging, the RIFE model [16] is employed to interpolate $16$ intermediate frames between adjacent sharp frames.
-Blurry frames from Cameras $04$, $10$, $16$, and $22$ are used for training, while the sharp frames from the remaining $18$ cameras are reserved for novel-view test.
-SMPL parameters are calibrated using the blurry video with EasyMoCap [7, 56], and human masks are obtained using SAM [25].
 Table 3: Quantitative ablation results on two datasets. We colorize results as best, second best, and third best.
@@
 1.
-SMPL parameters are calibrated using EasyMoCap [7], and human masks are generated using SAM [25].
+SMPL parameters are calibrated using EasyMoCap, and human masks are generated using SAM [24].
 Table 4: Quantitative ablation results for the inter-frame motion regularization loss $\mathcal{L}_{reg}$. We colorize better result of each sector.
@@
 Extensive evaluations demonstrate that the proposed model consistently outperforms existing baselines.
+Acknowledgment
+This research was supported in part by the Tateisi Science and Technology Foundation, JSPS KAKENHI Grant Numbers 24KK0209 and 22H00529, and Advanced AI Talent Development to Lead the Next-Generation AI for Intelligent Society (BOOST NAIS) of The University of Tokyo.